2. Income Statement Sales Costs Balance Sheet $25,400 Assets $61,000 Debt $26,900 Equity 34,100 17,300 Taxable 8,100 Total $61.000 Total $ 61000 income Taxes (21%) 1701 Net incomes 6,399 Assets and costs are proportional to sales. Debt and equity are not. A dividend of $2,100 was paid, and the company wishes to maintain a constant payout ratio. Next years sales are projected to be $29,210. What is the external financing needed? (Do not round intermediate calculations.)
Answer #1

Growth rate in sales=(29210-25400)/25400=15%

Dividend payout ratio=Dividend/Net income

=(2100/6399)=0.328176277

Sales 29210
Costs(17300*1.15) 19895
Taxable income $9315
Tax(21%*$9315) $1956.15
Net income $7358.85
Less:Dividends($7358.85*0.328176277) $2415
Addition to retained earnings $4943.85

Total assets would be=$61000*1.15=$70150

Total equity=$34100+Addition to retained earnings

=$34100+$4943.85

=$39043.85

Total assets=Total equity+Total liabilities

Hence external financing needed=70150-$39043.85-$26900

=$4206.15